# PDMLP: Patch-based Decomposed MLP for Long-Term Time Series Forecasting

### v2

@@ Abstract
Recent studies have attempted to refine the Transformer architecture to demonstrate its effectiveness in Long-Term Time Series Forecasting (LTSF) tasks. Despite surpassing many linear forecasting models with ever-improving performance, we remain skeptical of Transformers as a solution for LTSF. We attribute the effectiveness of these models largely to the adopted Patch mechanism, which enhances sequence locality to an extent yet fails to fully address the loss of temporal information inherent to the permutation-invariant self-attention mechanism. Further investigation suggests that simple linear layers augmented with the Patch mechanism may outperform complex Transformer-based LTSF models. Moreover, diverging from models that use channel independence, our research underscores the importance of cross-variable interactions in enhancing the performance of multivariate time series forecasting. The interaction information between variables is highly valuable but has been misapplied in past studies, leading to suboptimal cross-variable models. Based on these insights, we propose a novel and simple Patch-based Decomposed MLP (PDMLP) for LTSF tasks. Specifically, we employ simple moving averages to extract smooth components and noise-containing residuals from time series data, engaging in semantic information interchange through channel mixing and specializing in random noise with channel independence processing. The PDMLP model consistently achieves state-of-the-art results on several real-world datasets. We hope this surprising finding will spur new research directions in the LTSF field and pave the way for more efficient and concise solutions.
1 Introduction
@@
Figure 1: The self-attention scores of a 2-layer Transformer with different Patch size trained on ETTh1. We follow the setup of PathcTST Nie et al. (2022), retaining only the Encoder while replacing the Decoder with a simple MLP, and using a channel independent approach. A patch size of 1 is equivalent to the original Transformer, indicating that time series data often exhibits a trend of being segmented into patch Zhang & Yan (2022); Tang & Zhang (2023), and an increase in patch size can mitigate this to some extent.
@@
Long-term Time Series Forecasting (LTSF) is a critical area of research in the fields of statistics and machine learning, aimed at using historical data to predict the future of one or more variables over a certain period Oreshkin et al. (2019); Gong et al. (2023); Lin et al. (2023). Time series data are organized in chronological order and reveal underlying dynamic patterns, be they cyclical or non-cyclical Cryer (1986). Such forecasts play a vital role in various sectors, including biomedicine Liu et al. (2018), economics and finance Patton (2013), electricity Zhou et al. (2021), and transportation Yin & Shang (2016).
@@
Multivariate Time Series (MTS) consist of multiple variables recorded at the same time point, where each dimension may represent an individual univariate time series or be considered as a signal with multiple channels. With the advancement of deep learning, a number of models have been developed to enhance the performance of MTS forecasting Borovykh et al. (2017); Bai et al. (2018); Liu et al. (2020). In particular, recent models based on the Transformer Vaswani et al. (2017) architecture have demonstrated significant potential in capturing long-term dependencies Radford et al. (2018); Devlin et al. (2018). In recent years, Transformers have emerged as a leading architecture in time series forecasting Lim et al. (2021); Liu et al. (2022); Shen et al. (2023), initially applied in the field of Natural Language Processing (NLP) Radford et al. (2019); Brown et al. (2020) and subsequently expanded to other domains such as Computer Vision (CV) Liu et al. (2021b); Feichtenhofer et al. (2022) through the adaptation of the Patch method, becoming a versatile framework.
@@
Originally, these models utilized a channel mixing approach Li et al. (2023b), which involves projecting vectors from different channels recorded at the same time point into an embedding space and blending the information Zhou et al. (2021); Tang & Zhang (2023) ,and many models adopt the idea of time series decomposition Wu et al. (2021); Zhou et al. (2022); Wang et al. (2023). Nonetheless, recent studies have shown that a channel independence model might be more effective Nie et al. (2022); Han et al. (2023). Channel independence means that each input token contains information from only one channel, and intuitively, the models treat the MTS as separate univariate series for individual processing. These studies suggest that for LTSF tasks, an emphasis on channel independence may be more effective than channel mixing strategies Gong et al. (2023).
@@
Moreover, the effectiveness of Transformers in LTSF tasks has been challenged by a reconsideration of the Multi-Layer Perceptron (MLP) Zeng et al. (2023); Ekambaram et al. (2023); Chen et al. (2023), whose surprisingly simple architecture has outperformed all previous Transformer models. This raises a compelling question: Are Transformers effective for LTSF tasks? In response to this skepticism, recent Transformer-based models have adopted patch-based representations and achieved noteworthy performance in LTSF Nie et al. (2022); Chen et al. (2024).
@@
  • Does the impressive performance of the Patch-based Transformer derive from the inherent strength of the Transformer architecture, or is it merely due to the use of patch as the input representation?
@@
In this paper, we specifically discuss the three issues mentioned above and introduce the PDMLP, a novel and concise Patch-based decomposed MLP model tailored for LTSF task. PDMLP is entirely based on fully connected networks and incorporates the concept of patch. Specifically, we employ a patch methodology to embed the original sequences into representational spaces, then extract the smooth components and noisy residuals of the series using a simple moving average technique for separate processing. We independently process stochastic noise across channels and facilitate semantic information interchange between variables through channel mixing. Upon evaluation on numerous real-world datasets, the PDMLP model demonstrated state-of-the-art (SOTA) performance. To be precise, our contributions can be summarized in three aspects:
@@
  • We analyze the effectiveness of patch in time series forecasting and propose a Multi-Scale Patch Embedding (MPE) approach. Unlike previous embedding methods that used a single linear layer, MPE is capable of capturing multiscale relationships between input sequences.
@@
  • We introduce a novel entirely MLP-based model, named PDMLP. By utilizing moving averages, it performs the decomposition of latent vector and adopts a different approach to channel mixing for semantic information exchange across variables.
@@
  • Our experiments across a wide range of datasets in various fields show that PDMLP consistently achieves SOTA performance across multiple forecasting benchmarks. Moreover, we conduct an extensive analysis of Patch-based methods with the aim of charting new directions for future research in time series forecasting.
2 Patch for Long-Term Time Series Forecasting
@@
Why is Patch effective in time series forecasting? Lee et al. (2023); Zhong et al. (2023) We have studied a patch-based Transformer model adopting channel independent mechanisms, and we set the decoder layer as a simple single-layer MLP for time series forecastine Nie et al. (2022). As shown in the Figure  2(b), we can clearly see that for the same input length, as the size of the patch increases, the overall Mean Squared Error (MSE) of the model exhibits a trend of decreasing and then increasing, or decreasing and then stabilizing. As the input length increases, the patch size required for the model to achieve optimal performance also gradually increases. In the original input data, the effect of Attention is so poor that it cannot even outperform a simple single-layer MLP, which raises the question of whether Attention might not be the optimal choice for time series modeling to some extent, as it cannot handle long-term time series well.
@@
Compared to textual data in the NLP field, original time series data contains many redundant features due to high-frequency sampling, and is easily influenced by noise to some extent during the sampling process Tang & Zhang (2022). We believe that for data with excessive noise in the original data, the original Transformer’s Attention mechanism is not the optimal choice as it cannot effectively eliminate the noise. Sparse Attention performs better because this sparse mechanism can effectively mitigate the impact of noise to some extent Li et al. (2019); Wu et al. (2020); Liu et al. (2021a), but it may also attenuate the impact of original features.
@@
Patch, on the other hand, compresses the data, reduces the dimensionality of the input data, and decreases redundant features. Additionally, patch provides a certain degree of smoothing, which can reduce the influence of outliers to some extent, and help filter out fluctuations and random noise, retaining more stable and representative information. In the CV field, patch also works well in ViT Dosovitskiy et al. (2020), MAE He et al. (2022). Time series data usually contains patterns at different scales, and patch provides a modeling of short-term time series, enhancing local information in the sequence, allowing the model to better learn and capture local features. Therefore, we believe that the effectiveness of Transformer is not due to the effect of Attention, but rather due to the presence of patch.
@@
Figure 3: Overall structure of PDMLP. (a) Raw time series of different variates are independently embedded as tokens. (b)The Feature Decomposition utilizing moving averages to decompose the tokens into the smooth components and noisy residuals. (c) The MLP processed sequences in two ways: intra-variable and inter-variable. (d) A projection maps the latent vectors back to future sequences.
@@
So, is a larger patch always better? In Figure  2(b), a larger patch does not seem to achieve better results, which may be partly influenced by the size of the hidden layer ($d_{model}$). With a larger patch size, more time points are projected into a fixed-size $d_{model}$, which may lead to excessive compression. As shown in Figure  2(c), as $d_{model}$ increases, the performance of the model gradually improves, indicating that larger patch perform better. In the extreme case, a patch that treats the entire input sequence as one patch and projects it into a vector is similar to iTransformer Liu et al. (2023). However, a larger $d_{model}$ indicates that there are more parameters for the model to learn, which can easily lead to underfitting and result in decreased model performance.
3 PDMLP
@@
As shown in Figure  3, PDMLP consists of four network components: Multi-Scale Patch Embedding layer, Feature Decomposition layer, Multi-Layer Perceptron (MLP) layer, Projection layer. The Multi-Scale Patch Embedding layer embeds the multivariate time series into latent space. The Feature Decomposition layer decomposes the latent vector into the smooth components and noisy residuals, then operate separately through MLP layer. Finally, the latent vector are mapped back to the feature space through the Projection layer to obtain the future sequences $\hat{\mathcal{X}}$. Next, we will introduce the above modules separately.
3.1 Multi-Scale Patch Embedding
Time series analysis relies on the accurate identification of local information within the sequence and optimization of the model. The use of patch can provide the model with a local view of the time series in the short term, thereby enhancing the representation of local information within the sequence and enabling the model to learn and capture these local features more precisely. However, traditional methods typically employ patch of a single-scale to embed the raw time series, which makes the model inclined to learn single-scale temporal relationships while neglecting the multi-scale nature and complexity of time series data.
@@ 3.1 Multi-Scale Patch Embedding
To capture local information in more detail and to fully understand the temporal relationships within the time series, we adopt multi-scale patche. This includes shorter patch to capture local high-frequency patterns, as well as longer patch to unearth long-term seasonal, trends and periodic fluctuations. Through the introduction of multi-scale patch, the model can flexibly learn representative features over different lengths of time spans, thus enhancing predictive accuracy and model generalization capabilities.
To decompose the multivariate time series $\mathcal{X}$ into a univariate series $x$, we have developed a suite of patch $\mathcal{P}$ across various scales to process $x$. For a particular scale $p\in\mathcal{P}$, $x$ is first divided into non-overlapping patch where the length of patch corresponds to $p$, thus the patching process yields the sequence of patch $x_{p}\in\mathbb{R}^{N\times p}$, where $N$ is the number of patches. Subsequently, we employ a single-layer linear layer to embed the patch $x_{p}$, resulting in latent vectors $x_{e}\in\mathbb{R}^{N\times d}$, $d$ is the embedding dimension, the $d$ corresponding to different scales patch can be different. These latent vectors are then unfolded and concatenated with latent vectors obtained from patch at alternative scales, thus assembling the ultimate embedding vector $X\in\mathbb{R}^{1\times d_{model}}$, $d_{model}$ is the final embedding dimension input into the model. This multi-scale patch strategy permits the model to capture and learn the intrinsic dynamics of time series at different levels, thereby providing more precise insights when forecasting future trends and patterns.
3.2 Feature Decomposition
@@
Figure 4: Overall structure of MLP layer. (a) The embedded vectors first interact with the temporal information within the variable through the Intra-Variable MLP. (b) Then interact with the feature domain information between variables through the Intra-Variable MLP. (c) Subsequently, they are multiplied by the input of the Inter-Variable MLP using a dot-product approach. (d) Finally, they are added to the initial input of the MLP Layer using skip connections.
3.3 MLP layer
@@
Residual Connections: Residual connections He et al. (2016) are applied after each MLP, enabling the model to learn deeper architectures more efficiently. These connections not only help to mitigate the problems of vanishing gradients in deep networks but also provide the model with a shortcut path that ensures key temporal or feature information is not overlooked.
@@
While the minimalist architecture of PDMLP may not appear as eye-catching as some of the recently proposed, more complex Transformer models, empirical results indicate that PDMLP exhibits competitive performance, both in terms of training efficiency and accuracy on standard benchmark tests, relative to SOTA models.
@@
Table 1: Full results of the Long-Term Time Series Forecasting task. For all baselines, we adhere to the setting of the iTransformer with an input sequence length of 96. We conducted comparisons among an array of competitive models during various forecast horizons, and the outcomes were obtained from the iTransformer. Avg denotes the average result across all four prediction lengths. We denote the optimal results with boldface, and the second-best with an underline for clarity.

### v3

# Unlocking the Power of Patch: Patch-Based MLP for Long-Term Time Series Forecasting

@@ Abstract
Recent studies have attempted to refine the Transformer architecture to demonstrate its effectiveness in Long-Term Time Series Forecasting (LTSF) tasks. Despite surpassing many linear forecasting models with ever-improving performance, we remain skeptical of Transformers as a solution for LTSF. We attribute the effectiveness of these models largely to the adopted Patch mechanism, which enhances sequence locality to an extent yet fails to fully address the loss of temporal information inherent to the permutation-invariant self-attention mechanism. Further investigation suggests that simple linear layers augmented with the Patch mechanism may outperform complex Transformer-based LTSF models. Moreover, diverging from models that use channel independence, our research underscores the importance of cross-variable interactions in enhancing the performance of multivariate time series forecasting. The interaction information between variables is highly valuable but has been misapplied in past studies, leading to suboptimal cross-variable models. Based on these insights, we propose a novel and simple Patch-based MLP (PatchMLP) for LTSF tasks. Specifically, we employ simple moving averages to extract smooth components and noise-containing residuals from time series data, engaging in semantic information interchange through channel mixing and specializing in random noise with channel independence processing. The PatchMLP model consistently achieves state-of-the-art results on several real-world datasets. We hope this surprising finding will spur new research directions in the LTSF field and pave the way for more efficient and concise solutions. Code is available at: https://github.com/TangPeiwang/PatchMLP
Introduction
@@
Figure 1: The self-attention scores of a 2-layer Transformer with different Patch size trained on ETTh1. We follow the setup of PathcTST (Nie et al. 2022), retaining only the Encoder while replacing the Decoder with a simple MLP, and using a channel independent approach. A patch size of 1 is equivalent to the original Transformer, indicating that time series data often exhibits a trend of being segmented into patches (Zhang and Yan 2022; Tang and Zhang 2023), and an increase in patch size can mitigate this to some extent.
@@
Long-term Time Series Forecasting (LTSF) is a critical area of research in the fields of statistics and machine learning, aimed at using historical data to predict the future of one or more variables over a certain period (Oreshkin et al. 2019; Gong, Tang, and Liang 2023; Lin et al. 2023). Time series data are organized in chronological order and reveal underlying dynamic patterns, be they cyclical or non-cyclical (Cryer 1986). Such forecasts play a vital role in various sectors, including biomedicine (Liu et al. 2018), economics and finance (Patton 2013), electricity (Zhou et al. 2021), and transportation (Yin and Shang 2016).
@@
Multivariate Time Series (MTS) consist of multiple variables recorded at the same time point, where each dimension may represent an individual univariate time series or be considered as a signal with multiple channels. With the advancement of deep learning, a number of models have been developed to enhance the performance of MTS forecasting (Borovykh, Bohte, and Oosterlee 2017; Bai, Kolter, and Koltun 2018; Liu et al. 2020). In particular, recent models based on the Transformer (Vaswani et al. 2017) architecture have demonstrated significant potential in capturing long-term dependencies (Radford et al. 2018; Devlin et al. 2018). In recent years, Transformers have emerged as a leading architecture in time series forecasting (Lim et al. 2021; Liu et al. 2022b; Shen et al. 2023), initially applied in the field of Natural Language Processing (NLP) (Radford et al. 2019; Brown et al. 2020) and subsequently expanded to other domains such as Computer Vision (CV) (Liu et al. 2021b; Feichtenhofer et al. 2022) through the adaptation of the Patch method, becoming a versatile framework.
@@
Originally, these models utilized a channel mixing approach (Li et al. 2023b), which involves projecting vectors from different channels recorded at the same time point into an embedding space and blending the information (Zhou et al. 2021; Tang and Zhang 2023) ,and many models adopt the idea of time series decomposition (Wu et al. 2021; Zhou et al. 2022; Wang et al. 2023). Nonetheless, recent studies have shown that a channel independence model might be more effective (Nie et al. 2022; Han, Ye, and Zhan 2023). Channel independence means that each input token contains information from only one channel, and intuitively, the models treat the MTS as separate univariate series for individual processing. These studies suggest that for LTSF tasks, an emphasis on channel independence may be more effective than channel mixing strategies (Gong, Tang, and Liang 2023).
@@
Moreover, the effectiveness of Transformers in LTSF tasks has been challenged by a reconsideration of the Multi-Layer Perceptron (MLP) (Zeng et al. 2023; Ekambaram et al. 2023; Chen et al. 2023), whose surprisingly simple architecture has outperformed all previous Transformer models. This raises a compelling question: Are Transformers effective for LTSF tasks? In response to this skepticism, recent Transformer-based models have adopted patch-based representations and achieved noteworthy performance in LTSF (Nie et al. 2022; Chen et al. 2024).
@@
  • Does the impressive performance of the Patch-based Transformer derive from the inherent strength of the Transformer architecture, or is it merely due to the use of patches as the input representation?
@@
In this paper, we specifically discuss the three issues mentioned above and introduce the PatchMLP, a novel and concise Patch-based MLP model tailored for LTSF task. PatchMLP is entirely based on fully connected networks and incorporates the concept of patch. Specifically, we employ a patch methodology to embed the original sequences into representational spaces, then extract the smooth components and noisy residuals of the series using a simple moving average technique for separate processing. We independently process stochastic noise across channels and facilitate semantic information interchange between variables through channel mixing. Upon evaluation on numerous real-world datasets, the PatchMLP model demonstrated state-of-the-art (SOTA) performance. To be precise, our contributions can be summarized in three aspects:
@@
  • We analyze the effectiveness of Patches in time series forecasting and propose a Multi-Scale Patch Embedding (MPE) approach. Unlike previous embedding methods that used a single linear layer, MPE is capable of capturing multiscale relationships between input sequences.
@@
  • We introduce a novel entirely MLP-based model, named PatchMLP. By utilizing moving averages, it performs the decomposition of latent vector and adopts a different approach to channel mixing for semantic information exchange across variables.
@@
  • Our experiments across a wide range of datasets in various fields show that PatchMLP consistently achieves SOTA performance across multiple forecasting benchmarks. Moreover, we conduct an extensive analysis of Patch-based methods with the aim of charting new directions for future research in time series forecasting.
Patch for Long-Term Time Series Forecasting
@@
Why is Patch effective in time series forecasting? (Lee, Park, and Lee 2023; Zhong et al. 2023) We have studied a patch-based Transformer model adopting channel independent mechanisms, and we set the decoder layer as a simple single-layer MLP for time series forecastine (Nie et al. 2022). As shown in the Figure  2(b), we can clearly see that for the same input length, as the size of the patch increases, the overall Mean Squared Error (MSE) of the model exhibits a trend of decreasing and then increasing, or decreasing and then stabilizing. As the input length increases, the patch size required for the model to achieve optimal performance also gradually increases. In the original input data, the effect of Attention is so poor that it cannot even outperform a simple single-layer MLP, which raises the question of whether Attention might not be the optimal choice for time series modeling to some extent, as it cannot handle long-term time series well.
@@
Compared to textual data in the NLP field, original time series data contains many redundant features due to high-frequency sampling, and is easily influenced by noise to some extent during the sampling process (Tang and Zhang 2022). We believe that for data with excessive noise in the original data, the original Transformer’s Attention mechanism is not the optimal choice as it cannot effectively eliminate the noise. Sparse Attention performs better because this sparse mechanism can effectively mitigate the impact of noise to some extent (Li et al. 2019; Wu et al. 2020; Liu et al. 2021a), but it may also attenuate the impact of original features.
@@
Patch, on the other hand, compresses the data, reduces the dimensionality of the input data, and decreases redundant features. Additionally, patch provides a certain degree of smoothing, which can reduce the influence of outliers to some extent, and help filter out fluctuations and random noise, retaining more stable and representative information. In the CV field, patch also works well in ViT (Dosovitskiy et al. 2020), MAE (He et al. 2022). Time series data usually contains patterns at different scales, and patch provides a modeling of short-term time series, enhancing local information in the sequence, allowing the model to better learn and capture local features. Therefore, we believe that the effectiveness of Transformer is not due to the effect of Attention, but rather due to the presence of patch.
@@
Figure 3: Overall structure of PatchMLP. First, the raw time series of different variables are independently processed through Multi-scale Patch Embedding. Then, Feature Decomposition uses moving averages to decompose the embedded tokens into smooth components and noisy residues. Next, a MLP processes the sequences in two ways: intra-variable and inter-variable. Finally, the Predictor maps the latent vectors back to predictions and aggregates them into future series.
@@
So, is a larger patch always better? In Figure  2(b), a larger patch does not seem to achieve better results, which may be partly influenced by the size of the hidden layer ($d_{model}$). With a larger patch size, more time points are projected into a fixed-size $d_{model}$, which may lead to excessive compression. As shown in Figure  2(c), as $d_{model}$ increases, the performance of the model gradually improves, indicating that larger patches perform better. In the extreme case, a patch that treats the entire input sequence as one patch and projects it into a vector is similar to iTransformer (Liu et al. 2023). However, a larger $d_{model}$ indicates that there are more parameters for the model to learn, which can easily lead to underfitting and result in decreased model performance.
PatchMLP
@@
As shown in Figure  3, PatchMLP consists of four network components: Multi-Scale Patch Embedding layer, Feature Decomposition layer, Multi-Layer Perceptron (MLP) layer, Projection layer. The Multi-Scale Patch Embedding layer embeds the multivariate time series into latent space. The Feature Decomposition layer decomposes the latent vector into the smooth components and noisy residuals, then operate separately through MLP layer. Finally, the latent vector are mapped back to the feature space through the Projection layer to obtain the future sequences $\hat{\mathcal{X}}$. Next, we will introduce the above modules separately.
Multi-Scale Patch Embedding
Time series analysis relies on the accurate identification of local information within the sequence and optimization of the model. The use of patches can provide the model with a local view of the time series in the short term, thereby enhancing the representation of local information within the sequence and enabling the model to learn and capture these local features more precisely. However, traditional methods typically employ patches of a single-scale to embed the raw time series, which makes the model inclined to learn single-scale temporal relationships while neglecting the multi-scale nature and complexity of time series data.
@@
To capture local information in more detail and to fully understand the temporal relationships within the time series, we adopt multi-scale patch. This includes shorter patch to capture local high-frequency patterns, as well as longer patch to unearth long-term seasonal, trends and periodic fluctuations. Through the introduction of multi-scale patch, the model can flexibly learn representative features over different lengths of time spans, thus enhancing predictive accuracy and model generalization capabilities.
To decompose the multivariate time series $\mathcal{X}$ into a univariate series $x$, we have developed a suite of patches $\mathcal{P}$ across various scales to process $x$. For a particular scale $p\in\mathcal{P}$, $x$ is first divided into non-overlapping patches where the length of patch corresponds to $p$, thus the patching process yields the sequence of patches $x_{p}\in\mathbb{R}^{N\times p}$, where $N$ is the number of patches. Subsequently, we employ a single-layer linear layer to embed the patches $x_{p}$, resulting in latent vectors $x_{e}\in\mathbb{R}^{N\times d}$, $d$ is the embedding dimension, the $d$ corresponding to different scales patches can be different. These latent vectors are then unfolded to obtain the final embedding vectors $X\in\mathbb{R}^{1\times d_{model}}$, $d_{model}$ is the final embedding dimension input into the model. This multi-scale patch strategy permits the model to capture and learn the intrinsic dynamics of time series at different levels, thereby providing more precise insights when forecasting future trends and patterns.
Feature Decomposition
@@
Figure 4: Overall structure of MLP layer. The embedded vectors first interact with the temporal information within the variable through the Intra-Variable MLP. Then interact with the feature domain information between variables through the Intra-Variable MLP. Subsequently, they are multiplied by the input of the Inter-Variable MLP using a dot-product approach. Finally, they are added to the initial input of the MLP Layer using skip connections.
MLP layer
@@
Residual Connections: Residual connections (He et al. 2016) are applied after each MLP, enabling the model to learn deeper architectures more efficiently. These connections not only help to mitigate the problems of vanishing gradients in deep networks but also provide the model with a shortcut path that ensures key temporal or feature information is not overlooked.
@@
While the minimalist architecture of PatchMLP may not appear as eye-catching as some of the recently proposed, more complex Transformer models, empirical results indicate that PatchMLP exhibits competitive performance, both in terms of training efficiency and accuracy on standard benchmark tests, relative to SOTA models.
Loss Function
Our loss function is calculated by the Mean Square Error (MSE) between the model prediction $\hat{x}$ and the ground truth $x$: $\mathcal{L}=\frac{1}{M}\sum_{M}^{i=1}\left\|\hat{x}^{i}_{L+1:L+T}-x^{i}_{L+1:L+T}\right\|^{2}_{2}$ and the loss is propagated back from the Projection output across the entire model.
Experiments
We comprehensively evaluated the proposed PatchMLP in various time series forecasting applications, verifying the generality of the proposed framework, and have further delved into the investigation of the effectiveness of individual components of PatchMLP when applied in LSTF tasks.
Datasets: We evaluated the performance of our proposed PatchMLP on 8 commonly used LSTF benchmark datasets: Solar Energy (Lai et al. 2018), Weather, Traffic, Electricity (ECL) (Wu et al. 2021), and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al. 2021).
@@
Table 1: Results of the Long-Term Time Series Forecasting task. For all baselines, we adhere to the setting of the iTransformer with an input sequence length of 96. We conducted comparisons among an array of competitive models during various forecast horizons. All the results are averaged from 4 different prediction lengths, that is {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction, we denote the optimal results with boldface for clarity.